Physical Chemistry

Number of coulomb required to deposit 5.4 grams of aluminium when the given electrode equation is Al³+ +3e---> Al

sreyaaw tallapragada
8 Years agoGrade

1 Answer

Arun
8 Years ago
Dear student
 
For one mole of Al (27 gm) requires = 3*96500 C
1 gm of Al requires = 3 *96500/27
Hence
5.4 gm requires = 3 * 96500* 5.4/27 
 = 57900 C
 
Regards
Arun (askIITians forum expert)